# NeSyGeo: A Neuro-Symbolic Framework for Multimodal Geometric Reasoning Data Generation

### v1

@@ Abstract
Obtaining large-scale, high-quality data with reasoning paths is crucial for improving the geometric reasoning capabilities of multi-modal large language models (MLLMs). However, existing data generation methods, whether based on predefined templates or constrained symbolic provers, inevitably face diversity and numerical generalization limitations. To address these limitations, we propose NeSyGeo, a novel neuro-symbolic framework for generating geometric reasoning data. First, we propose a domain-specific language grounded in the entity–relation–constraint paradigm to comprehensively represent all components of plane geometry, along with generative actions defined within this symbolic space. We then design a symbolic–visual–text pipeline that synthesizes symbolic sequences, maps them to corresponding visual and textual representations, and generates diverse question–answer (Q&A) pairs using large language models (LLMs). To the best of our knowledge, we are the first to propose a neuro-symbolic approach in generating multimodal reasoning data. Based on this framework, we construct NeSyGeo-CoT and NeSyGeo-Caption datasets, containing 100k samples, and release a new benchmark NeSyGeo-Test for evaluating geometric reasoning abilities in MLLMs. Experiments demonstrate that the proposal significantly and consistently improves the performance of multiple MLLMs under both reinforcement and supervised fine-tuning. With only 4k samples and two epochs of reinforcement fine-tuning, base models achieve improvements of up to +15.8% on MathVision, +8.4% on MathVerse, and +7.3% on GeoQA. Notably, a 4B model can be improved to outperform an 8B model from the same series on geometric reasoning tasks.
@@ 1 Introduction
Improving the visual reasoning capabilities of MLLMs has garnered significant attention recently  [17, 2, 1, 14, 11, 33, 25, 28, 15], with models like InternVL [5] and the QwenVL series [27, 3] demonstrating significant enhancements in visual-semantic comprehension through their multimodal capabilities. Among various visual reasoning tasks, geometric mathematical reasoning is crucial for evaluating the reasoning performance of MLLMs [31, 30], as it requires a deep integration of spatial perception, symbolic understanding, and logical deduction. To enhance such reasoning abilities, existing approaches [37, 38, 39] primarily rely on fine-tuning base models using reinforcement learning (RL) or supervised fine-tuning (SFT) on specialized geometric reasoning datasets. These methods depend heavily on the availability of large-scale, high-quality geometric reasoning data, which is often costly and time-consuming to construct manually. Therefore, automatic data generation for geometric reasoning has emerged as a promising and actively explored direction, aiming to alleviate data scarcity and further improve the reasoning abilities of MLLMs.
Existing approaches for generating datasets in geometric tasks can be broadly classified into four categories. Text augmentation methods like G-LLaVA [9] primarily mutate the conditions of
@@
existing datasets through equivalent condition transformation and numerical scaling. However, this approach fails to address the scalability of image generation. Template-based methods [7, 37, 12], use predefined geometric templates with fixed topologies, simplifying synthesis but constraining diversity by reducing the geometric space to limited combinations. Solver-based methods [10, 32] inspired by symbolic prover AlphaGeometry [23], leverage formal languages for synthesis but lack metric details (e.g., angles, lengths, areas), restricting multimodal data to descriptive annotations and limiting numerical reasoning applications. Tool-based methods attempt to generate codes for tools like GeoGebra or MATLAB via LLMs. However, even advanced models struggle to ensure correctness with ambiguous natural language instructions and complex geometric spaces. In summary, existing methods grapple with issues of image scalability, limited geometric diversity, a lack of precise numerical information, and challenges in ensuring the reliability of generated content.
Beyond the challenges in data synthesis methodologies, current geometric reasoning datasets present several limitations that impede the advancement of MLLMs. A primary limitation stems from the often inadequate quality and low resolution of the provided images. Such inputs frequently fall below the optimal requirements of visual encoders [20, 16, 17], hindering the extraction of crucial fine-grained visual features and discriminative information essential for robust multimodal reasoning. Furthermore, our analysis reveals notable information redundancy between the textual and visual modalities in many current datasets. As shown in Figure 1, our comparative experiments demonstrate minimal or negligible accuracy drops upon image removal. This finding emphasizes the urgent need for a dataset that effectively separates textual and visual information and provides high-quality images to promote MLLMs’ visual perception and logical reasoning performance.
@@
To address these challenges, we propose NeSyGeo, a neuro-symbolic framework for synthesising high-quality multimodal geometric reasoning datasets. NeSyGeo integrates three components: 1) A formal geometric symbolic space defined by a domain-specific language (DSL), capturing primitive entities (points, lines, circles), topological relations (parallelism, incidence, perpendicularity), and metric constraints (angles, lengths), enabling diverse geometric configurations via systematic sampling within constrained parametric bounds. 2) A bidirectional conversion engine that transforms symbolic constructs into decoupled modalities, producing annotated vector graphics paired with concise textual axioms. 3) A causal Q&A pairs and theorem-grounded Chain-of-Thought (CoT) sequences generator that effectively merges neural reasoning with symbolic verification. To our knowledge, we are the first to develop a neuro-symbolic framework for producing multimodal reasoning data.
Our framework enhances the diversity and validity of generated geometric reasoning data while effectively mitigating information redundancy and the underutilization of visual signals during training. Specifically, the comprehensive Geo-DSL and its expansive symbolic synthesis action-space promote diverse and well-grounded image generation. Meanwhile, our CoT sequence generator, powered by LLMs’ strong reasoning and language capabilities, conducts a backwards search across the geometric space to construct Q&A pairs, thereby enriching textual diversity. The unique identification of geometric elements via our symbolic language and dedicated conversion engine ensures visual validity. In parallel, a bidirectional cross-validation process using expert LLMs ensures textual validity. By strategically distributing complementary information across image and text modalities, our approach
@@
Figure 2: Comparison of dataset characteristics synthesized by our method and other popular synthesis approaches. “High Resolution” denotes average image pixels exceeding 336$\times$336. “Symbolic Form” refers to the symbolic meta-information associated with the image. “Classification of Elements” signifies categorization by geometric elements. “Visual Understanding” represents the mitigation of image-text redundancy for stronger visual grounding in reasoning. More specific examples of different methods are in Appendix A.
encourages MLLMs to actively engage with visual information when solving problems, enhancing their abilities to perceive visually and effectively utilize images.
Leveraging the NeSyGeo pipeline, we construct two training datasets, NeSyGeo-Caption and NeSyGeo-CoT, comprising 100k samples. NeSyGeo-Caption aims to improve the perceptual understanding of geometric elements, while NeSyGeo-CoT primarily focuses on enhancing logical reasoning. The key characteristics of our dataset compared to other popular multimodal geometric datasets are presented in Figure 2. Additionally, we develop an evaluation set, NeSyGeo-Test, with 2668 Q&A pairs, enabling a thorough assessment of the geometric reasoning capabilities of mainstream MLLMs. We conducted an extensive and comprehensive evaluation of the geometric reasoning capabilities of current mainstream open-source and closed-source models, with details presented in Appendix 8. Notably, our training dataset consistently and efficiently enhances the geometric reasoning performance of MLLMs across multiple benchmarks. With only 4k samples and two epochs of RL training, base models achieve performance improvements of up to +15.8% on MathVision, +8.4% on MathVerse, and +7.3% on GeoQA. Moreover, InternVL2.5-4B can be improved to outperform the 8B model in the same series on geometric reasoning tasks.
@@
  • We propose NeSyGeo, a novel framework for geometric reasoning data generation, featuring a Geo-DSL for symbolic synthesis, a conversion engine for image and text generation, and an LLM-driven generator for Q&A pairs with CoT. NeSyGeo ensures validity through rigorous symbolic definitions and diversity via varied actions and neural searching.
  • Using our framework, we synthesize the NeSyGeo-Caption and NeSyGeo-CoT training datasets with 100k high-quality samples, alongside a comprehensive geometric task evaluation set NeSyGeo-Test. These datasets are characterized by their diversity, rigor, and balanced distribution of information across image and text modalities.
@@ 2.1 Geometric Problem-Solving
Early approaches to geometric reasoning predominantly relied on symbolic solvers that used formal languages to tackle the tasks. For instance, Inter-GPS [18] and PGDP [35] employed symbolic methods by manually crafting reasoning rules and symbolic representations for geometric entities. These systems typically transform visual input into symbolic forms through instance segmentation and apply theorem search to derive solutions. However, these methods lacked scalability due to their dependence on manually designed rules. Their inability to generalize beyond specific problem types further limited their universality and effectiveness across diverse geometric challenges.
Figure 3: The overview of our neuro-symbolic data generation framework. The framework comprises three steps: In the first step, we get a symbolic language sequence in a limited symbolic action space. In the second step, the conversion engine parses the Geo-DSL sequence and translates it back to natural language and visual image without losing soundness. In the third step, we employ LLMs to take a reverse search and forward validation process to get final Q&A pairs with CoT.
The advent of MLLMs has shifted the paradigm toward data-driven geometric reasoning, leveraging their robust reasoning capabilities. Recent advancements include GeoDRL [19]and GeoGen [13]. Despite these developments, geometric reasoning poses significant challenges for MLLMs, requiring seamless integration of image perception, geometric knowledge, and multi-step reasoning. GeoSense [29] identifies the identification and application of geometric principles as a persistent bottleneck. Similarly, GeoEval [34] reveals that current MLLMs exhibit significantly low accuracy when facing more challenging geometric problems. MathVerse [36] further highlights MLLMs’ over-reliance on textual information, underscoring the critical need for balanced multimodal datasets to enhance cross-modal reasoning capabilities.
@@ 2.2 Multimodal Geometry Datasets
Large-scale, high-quality datasets are essential for enhancing the performance of MLLMs in solving geometric problems. Early datasets such as GeoS [21] (186 problems), Geometry3k [18] (3000 problems) and GeoQA [4] (4998 problems) utilized human manual annotation. Their datasets are thus limited to a small scale. With the development of MLLMs, datasets of greater magnitude have become essential. To address this, numerous efforts have shifted toward automatic data generation.
G-LLaVA [9] rephrased questions from GeoQA and Geometry3k to create 115,000 Q&A pairs, but failed to enhance image variety. Template-based methods [37, 7] typically rely on 10–20 predefined geometric figures, limiting the diversity of the generated images. AlphaGeometry [23], a notable work that combines symbolic solvers for geometric proofs, employs a symbolic language definition. Yet, due to the absence of numerical attributes such as angle measures and segment lengths in its geometric space, attempts to automatically generate datasets using the AlphaGeometry framework [10, 32] are confined to caption datasets, failing to produce the numerical Q&A pairs critical for current MLLMs training. In contrast to prior approaches, our method pioneers a neuro-symbolic framework, being the first to integrate the precision of symbolic definition with the diversity of neural search for generating multimodal reasoning data.
@@ 3 Methods
To address the urgent need for large-scale and high-quality multimodal datasets in MLLMs for geometric reasoning, we propose NeSyGeo, a novel three-stage data generation pipeline. The pipeline is built upon Geo-DSL, a symbolic DSL designed to represent most elements in plane geometry space concisely and wholly. Its entity-relation-constraint structure allows any element to
Table 1: Examples of our Geo-DSL and Corresponding Natural Language. Geo-DSL is defined by an entity-relation-constraint framework, encompassing 13 point types, 7 line types, 3 angle types, and 14 shape types in plane geometry. See Appendix H for complete Geo-DSL definitions.
Type Geo-DSL Language Natural Language
Shape $Triangle(A,B,C)=(x,y,\alpha)$ Triangle $ABC$ has $AB=x$, $BC=y$, $\angle B=\alpha$
$Circle(O)=(x)$ Circle $O$ has radius $x$
Point $Foot(D,A,Line(B,C))$ $D$ is the foot of the perpendicular from $A$ to $BC$
$Intersection(E,Line(A,B),Line(C,D))$ $E$ is the intersection of line $AB$ and line $CD$
Line $Para(Line(A,B),Line(C,D),x)$ Line $AB$ is parallel to $CD$, $AB=x$
Angle $Angle(P,Q,R)=\alpha$ $\angle PQR=\alpha$
be defined via a single statement, while its expressive power ensures comprehensive coverage of all geometric elements and values.
NeSyGeo’s generation process unfolds in three distinct stages: First, a symbolic generator performs action augmentations within the finite symbolic space to produce a Geo-DSL sequence. This approach effectively constrains the synthesis and augmentation process, ensuring validity and controllability by generating outside the infinite domains of natural language and image spaces (Section 3.2). Second, a conversion engine maps the generated Geo-DSL sequences back into natural language descriptions and visual image representations. This process synthesizes high-quality images and valid text while avoiding intermodal information overlap. Third, to get Q&A Pairs with reasoning paths, we utilize expert LLMs to conduct backwards search to identify the geometric unknowns to be solved and generate the CoT in a forward manner (Section 3.4). The search process primarily ensures the diversity of the Q&A pairs, while the forward verification confirms the correctness of the CoT and the final answer. The overall framework of NeSyGeo is illustrated in Figure 3.
@@ 3.1 Symbolic Definition
Existing symbolic languages for plane geometry have certain limitations. The definitions in AlphaGeometry [23] are tailored for proof-based problems, thus lacking definitions related to specific numerical values. InterGPS [18] defines a predicate as a geometric shape entity, geometric relation, or arithmetic function, constructing 91 predicates. However, this approach overly fragments shape, attribute, and relation definitions into independent statements, often requiring multiple statements to specify a single element in the figure. This significantly increases the complexity of a conversion engine’s identification of elements within the symbolic space and further conversion of them.
We propose Geo-DSL, a concise and comprehensive symbolic language for plane geometry to address these limitations. It employs an entity-relation-constraint framework, using well-defined rules to uniquely define 13 types of points, 7 types of lines, 3 types of angles, and 14 types of shapes, covering all plane geometry elements and incorporating numerical attributes like lengths and angles for precise specifications. Table 1 provides partial examples of symbolic definitions. Geo-DSL offers two key advantages. First, its comprehensive coverage includes all geometric elements and their numerical properties, enabling accurate and complete descriptions. Second, its simplicity allows a single statement to specify an element uniquely, promoting the incremental integration of the symbolic action space and the sequential parsing of statements by the conversion engine. This combination of completeness and simplicity makes Geo-DSL an efficient and powerful geometric representation and processing solution.

### v2

@@ Abstract
Obtaining large-scale, high-quality reasoning data is crucial for improving the geometric reasoning capabilities of multi-modal large language models (MLLMs). However, existing data generation methods, whether based on predefined templates or constrained symbolic provers, inevitably face diversity and numerical generalization limitations. To address these limitations, we propose NeSyGeo, a novel neuro-symbolic framework for generating geometric reasoning data. First, we propose a domain-specific language grounded in the entity–attributes-relations paradigm to comprehensively represent all components of plane geometry, along with generative actions defined within this symbolic space. We then design a symbolic–visual–text pipeline that synthesizes symbolic sequences, maps them to visual and textual representations and generates reasoning path with reverse search and forward validation. Based on this framework, we construct NeSyGeo-CoT and NeSyGeo-Caption datasets, containing 100k samples, and release a new benchmark NeSyGeo-Test for evaluating geometric reasoning abilities in MLLMs. Experiments demonstrate that the proposal significantly and consistently improves the performance of multiple MLLMs under both reinforcement and supervised fine-tuning. With only 4k samples and two epochs of reinforcement fine-tuning, base models achieve improvements of up to +15.8% on MathVision, +8.4% on MathVerse, and +7.3% on GeoQA. Notably, a 4B model can be improved to outperform an 8B model from the same series on geometric reasoning tasks.
@@ 1 Introduction
Improving the visual reasoning capabilities of MLLMs has garnered significant attention recently  (Liu et al., 2023; Alayrac et al., 2022; Achiam et al., 2023; Li et al., 2021; Jiang et al., 2024; Zhang et al., 2025a; Wang et al., 2024a; Wu et al., 2024; Liang et al., 2024), with models like InternVL (Chen et al., 2024) and the QwenVL series (Wang et al., 2024c; Bai et al., 2025) demonstrating significant enhancements in visual-semantic comprehension through their multimodal capabilities. Among various visual reasoning tasks, geometric mathematical reasoning is crucial for evaluating the reasoning performance of MLLMs (Yan et al., 2025; 2024), as it requires a deep integration of spatial perception, symbolic understanding, and logical deduction. To enhance such reasoning abilities, existing approaches (Zhang et al., 2024d; 2025b; Zhao et al., 2025) primarily rely on fine-tuning base models using reinforcement learning (RL) or supervised fine-tuning (SFT) on specialized geometric reasoning datasets. These methods depend heavily on the availability of large-scale, high-quality geometric reasoning data, which is often costly and time-consuming to construct manually. Therefore, automatic data generation for geometric reasoning has emerged as a promising and actively explored direction, aiming to alleviate data scarcity and further improve the reasoning abilities of MLLMs.
@@
Existing approaches for generating datasets in geometric tasks can be broadly classified into four categories. Text augmentation methods like G-LLaVA (Gao et al., 2025) primarily mutate the conditions of existing datasets through equivalent condition transformation and numerical scaling. However, this approach fails to address the scalability of image generation. Template-based methods (Deng et al., 2024; Zhang et al., 2024d; Kazemi et al., 2024), use predefined geometric templates with fixed topologies, simplifying synthesis but constraining diversity by reducing the geometric space to limited combinations. Solver-based methods (Huang et al., 2025; Zhang et al., 2024a) inspired by symbolic prover AlphaGeometry (Trinh et al., 2024), leverage formal languages for synthesis but lack metric details (e.g., angles, lengths, areas), restricting multimodal data to descriptive annotations and limiting numerical reasoning applications.In summary, existing methods grapple with issues of image scalability, limited geometric diversity, a lack of precise numerical information, and challenges in ensuring the reliability of generated content.
Moreover, current geometric reasoning datasets present three critical limitations in cross-modal reasoning:1) Information Redundancy: Repetitive textual conditions provide a “shortcuts” that allows models to bypass visual reasoning. 2) Semantic Gap Between Modalities: Abstract geometric properties stated in the text (such as equality) are not explicitly grounded from the visual features. 3) Poor Image Quality: Low-resolution and inadequate images hinder the extraction of crucial visual information. As shown in Figure 1, our comparative experiments demonstrate minimal or negligible accuracy drops upon image removal. Such discrepancies hinder the model’s ability to make use of visual perception and learn robust cross-modal reasoning .
@@
To address these challenges, we propose NeSyGeo, a neuro-symbolic framework for synthesising high-quality multimodal geometric reasoning datasets. NeSyGeo integrates three components: 1) A formal geometric symbolic space defined by a domain-specific language (DSL), capturing primitive entities (points, lines, circles), metric attributes (angles, lengths) and topological relations (parallelism, incidence, perpendicularity), enabling diverse geometric configurations via systematic sampling within constrained parametric bounds. 2) Bidirectional conversion engines that transform symbolic constructs into decoupled modalities, producing annotated vector graphics paired with concise textual axioms. 3) Validated Question-Answer (Q&A) pairs and theorem-grounded CoT generators that effectively merges neural reasoning with verification.
Our framework enhances data quality by generating a wide spectrum of well-grounded images and correspondingly complex textual descriptions, while ensuring the validity and fidelity of both modalities. Furthermore, by carefully annotating semantic relationships within the image and avoiding redundancy between textual and visual conditions, it compels MLLMs to actively engage with visual information, thereby preventing modality neglect and significantly improving their visual grounding capabilities during training.
[…]
@@
Figure 2: Comparison of dataset characteristics synthesized by our method and other popular synthesis approaches. “High Resolution” denotes average image pixels exceeding $\times336$. “Symbolic Form” refers to the symbolic meta-information associated with the image. “Classification of Elements” signifies categorization by geometric elements. “Visual Understanding” represents the mitigation of image-text redundancy for stronger visual grounding in reasoning. More specific examples of different methods are in Appendix E.1.
@@
  • We propose NeSyGeo, a novel neuro-symbolic framework for geometric reasoning data generation. NeSyGeo incorporates a comprehensive Geo-DSL, a reasoner–verifier for CoT generation, and a painter–translator module for informalization. It ensures data quality, enhances diversity, improve the correctness of reasoning chains, and reduces redundancy across modalities.
  • Using our framework, we synthesize NeSyGeo-Caption and NeSyGeo-CoT training datasets with 100k samples, featuring high-quality images and reasoning chains, alongside a comprehensive geometric task evaluation set NeSyGeo-Test with 2,668 samples. These datasets are characterized by their diversity, rigor, and balanced distribution of information across image and text modalities.
@@ 2.1 Geometric Problem-Solving
Early approaches to geometric reasoning predominantly relied on symbolic solvers that used formal languages to tackle the tasks. For instance, Inter-GPS (Lu et al., 2021) and PGDP (Zhang et al., 2022) employed symbolic methods by manually crafting reasoning rules and symbolic representations for geometric entities. These systems typically transform visual input into symbolic forms through instance segmentation and apply theorem search to derive solutions. However, these methods lacked scalability due to their dependence on manually designed rules. Their inability to generalize beyond specific problem types further limited their universality across diverse geometric challenges.
The advent of MLLMs has shifted the paradigm toward data-driven geometric reasoning, leveraging their robust reasoning capabilities. Recent advancements include GeoDRL (Peng et al., 2023) and GeoGen (Krueger et al., 2021). Despite these developments, geometric reasoning poses significant challenges for MLLMs, requiring image perception, geometric knowledge, and multi-step reasoning. GeoSense (Xu et al., 2025) identifies the identification and application of geometric principles as a persistent bottleneck. Similarly, GeoEval (Zhang et al., 2024b) reveals that current MLLMs exhibit significantly low accuracy when facing more challenging geometric problems. MathVerse (Zhang et al., 2024c) further highlights MLLMs’ over-reliance on textual information, underscoring the critical need for balanced multimodal datasets to enhance cross-modal reasoning capabilities.
@@ 2.2 Multimodal Geometry Datasets
Large-scale, high-quality datasets are essential for enhancing the performance of MLLMs in solving geometric problems. Early datasets such as GeoS (Seo et al., 2015) (186 problems), Geometry3k (Lu et al., 2021) (3000 problems) and GeoQA (Chen et al., 2021) (4998 problems) utilized human manual annotation. Their datasets are thus limited to a small scale. With the development of MLLMs, datasets of greater magnitude have become essential. To address this, numerous efforts have shifted toward automatic data generation.
Figure 3: Our pipeline is centered around a symbolic language Geo-DSL. The Generator synthesizes a sequence in this language, from which Reasoner and Verifier produce logically sound Q&A pairs and reasoning chains. Subsequently, Painter and Translator render the symbolic core into semantically orthogonal visual and textual outputs.
G-LLaVA (Gao et al., 2025) rephrased questions from GeoQA and Geometry3k to create 115,000 Q&A pairs, but failed to enhance image variety. Template-based methods (Zhang et al., 2024d; Deng et al., 2024) typically rely on 10–20 predefined geometric figures, limiting the diversity of the generated images. AlphaGeometry (Trinh et al., 2024), a notable work that combines symbolic solvers for geometric proofs, employs a symbolic language definition. Yet, due to the absence of numerical attributes such as angle measures and segment lengths in its geometric space, attempts to automatically generate datasets using the AlphaGeometry framework.On the other hand,  (Huang et al., 2025) and  (Zhang et al., 2024a) are confined to caption datasets, failing to produce the numerical Q&A pairs critical for current MLLMs training. In contrast to prior approaches, our method pioneers a neuro-symbolic framework, being the first to integrate the precision of symbolic definition with the diversity of neural search for generating multimodal reasoning data.
@@ 3 Methods
The overall framework of NeSyGeo is illustrated in Figure 3 .The pipeline is built upon Geo-DSL(Section 3.1), with generation process unfolds in three parts:
First, a symbolic generator performs action augmentations within the finite symbolic space to produce a Geo-DSL sequence. This approach effectively constrains the synthesis and augmentation process, ensuring validity and controllability by generating outside the infinite domains of natural language and image spaces (Section 3.2).
Second, to get questions with reasoning paths, we utilize Reasoner and Verifier with a two-stage generation process. The reverse search primarily ensures the diversity of the Q&A pairs, while the forward verification confirms the correctness of the CoT and the final answer. (Section 3.3)
Third, painter and translator map the generated Geo-DSL sequences back into natural language descriptions and visual image respectively. This process synthesizes high-quality images and valid text while avoiding intermodal information overlap. (Section 3.4)
@@ 3.1 Symbolic Definition
Existing symbolic languages present significant limitations. The syntax of AlphaGeometry (Trinh et al., 2024), for example, is tailored for formal proofs and omits numerical specifications. In contrast, InterGPS (Lu et al., 2021) suffers from excessive fragmentation, requiring multiple statements to define single elements, which substantially complicates the parsing and conversion process.
To surmount these challenges, we introduce Geo-DSL, a concise and comprehensive symbolic language for plane geometry. It is architected around an entity–attributes-relations framework, which we denote as a triplet $\mathcal{G}=(\mathcal{E},\mathcal{A},\mathcal{R})$, representing entities, attributes, and relations respectively. The syntax supports three primary declarative constructs: (i) Attribute Assertions ($e\rightarrow A_{e}$), which bind quantitative properties to an entity; (ii) Relational Assertions ($(e_{1},\dots,e_{n})\rightarrow r$), which define relationships among entities; and (iii) Hybrid Declarations that compose the two.
Geo-DSL offers two key advantages. First, it achieves descriptive completeness through a structured ontology that classifies all geometric primitives into four fundamental types (points, lines, angles, and shapes), representing the plane geometry with only 37 core statements. Second, its syntactic parsimony, where each entity is unambiguously instantiated via a single statement, enables both the principled composition of the symbolic action space and its deterministic parsing by our conversion engine. This synthesis of expressive power and structural minimalism renders Geo-DSL a robust and scalable framework for automated geometric representation and processing.